The rotor is an amusement park ride where people stand against the inside of a cylinder. Once the cylinder is spinning fast enough, the floor drops out.

If the coefficient of static friction between a person and the wall of the cylinder is 0.600 and the cylinder has a radius of 3.80 m, what is the minimum angular speed of the cylinder so that the people don’t fall out?
1
Expert's answer
2020-10-27T11:22:48-0400

Given:-

Coefficient of static friction "\mu=0.600"

Radius of cylinder r=3.80m


Let "\omega" be the angular speed of rotor


To avoid slipping the frictional force must be balanced by the weight of the person

"\therefore mg=\mu\dfrac{mv^2}{r}\\\to g=\mu\dfrac{r^2\omega^2}{r}" ( as v=r"\omega)"


"\to \omega^2=\dfrac{g}{\mu r}"


"\to w=\sqrt{\dfrac{9.8}{0.600\times3.80}}"


"\to \omega=\sqrt{\dfrac{9.8}{2.28}}"


"\to \omega=\sqrt{4.71}=2.071rad s^{-1}"